A platform diver has a kinetic energy of 12,000 n x m prior to landing in the water bucket. If her mass is 40 kg, what is her final velocity before splash?


1
Expert's answer
2022-04-15T16:40:34-0400

"v=\sqrt\frac{2E}m=24.5~\frac ms."